OpenStudy (anonymous):

A)Solve 4x+x(x-1)=0 B)Find the x-intercepts of f(x)=4x+x(x-1) What are the solutions? What are the x-intercepts?

OpenStudy (cwtan):

4x+x^2-x=0 x^2+3x=0 x(x+3)=0 x=0 or 3 for x-intercept, f(x)=0 x=0 or 3

OpenStudy (anonymous):

So the x-intercept is x=0 or 3?

OpenStudy (cwtan):

yes

OpenStudy (anonymous):

What about the solution?

OpenStudy (cwtan):

The x-intercepts are the solution for (A)

OpenStudy (anonymous):

Ok thanks!!